# Causally Reliable Concept Bottleneck Models

### v2

@@ 5 Experimental evaluations
The considered datasets include both synthetic and real-world benchmarks. As synthetic datasets, we sample $10^{4}$ points from each of the five following discrete Bayesian networks available from the bnlearn repository (Scutari, 2010): Asia Lauritzen & Spiegelhalter (1988), Sachs Sachs et al. (2005), Insurance Binder et al. (1997), Alarm Beinlich et al. (1989) and Hailfinder Abramson et al. (1996). We include cMNIST, a variant of the original dataset (LeCun et al., 2010) in which the image data are colored according to custom rules. Additionally, we consider three real-world datasets: CelebA (Liu et al., 2015), a facial recognition dataset labeled with different binary facial attributes; CUBC, a custom version of the original bird image dataset (He & Peng, 2019) from which we select a subset of concepts and define new ones to introduce deeper causal relationships; Siim-Pneumothorax (You et al., 2023), containing chest X-ray images annotated with a single label indicating the presence of pneumothorax, without additional concepts or their annotations. To generate them, we follow the label-free approach outlined in Sec. 4.2.1. Exhaustive details on all datasets are given in App. E.
The performance of the proposed pipeline is investigated alongside an opaque neural baseline predicting the task variable only (OpaqNN) and established state-of-the-art (SOTA) concept-based architectures, namely: CBM (Koh et al., 2020), with linear and non-linear decoder; CEM (Zarlenga et al., 2022); and SCBM (Vandenhirtz et al., 2024). Hyperparameters have been selected via an independent search for each dataset–model pair based on performance on the validation set. Further details on each model’s architecture and hyperparameters can be found in App. F. Note that all baselines, except for OpaqNN, provide concept-based explanations for their predictions and allow concept interventions at test-time. This excludes architectures such as Self-Explainable Neural Networks (Alvarez Melis & Jaakkola, 2018) and Concept Whitening (Chen et al., 2020) as they do not offer a clear mechanism for intervening on their concept bottlenecks. We also excluded other CBM baselines such as Probabilistic CBMs (Kim et al., 2023), Post-hoc CBMs (Yuksekgonul et al., 2022), Label-free CBMs (Oikarinen et al., 2023; Yang et al., 2023), as all of them share the same limitation of vanilla CBMs and CEMs: the causal graph is fixed and bipartite. Python code for reproducing all experiments is provided alongside the submission as supplementary material.
@@ 5.1 Task accuracy and causal reliability
Table 1: Task accuracy (%). Task concepts are as follows: dysp (Asia), Akt (Sachs), PropCost (Insurance), BP (Alarm), R5Fcst (Hailfinder), parity (cMNIST), mouth slightly open (CelebA), Survival (CUBC), pneumothorax (Pneumoth.). refers to reduced concept bottlenecks. Methods with significant improvement are in bold. Uncertainties represent $2$ sample mean $\sigma$ (5 seeds).
@@
CBM+mlp $71.2_{\pm 1.4}$ $65.68_{\pm.84}$ $66.7_{\pm 1.4}$ $62.3_{\pm 1.8}$ $70.9_{\pm 2.2}$ $93.55_{\pm.31}$ $71.27_{\pm.20}$ $56.2_{\pm 1.9}$ $76.7_{\pm 0.6}$ $58.6_{\pm 2.7}$ $52.8_{\pm 1.2}$
CEM $71.1_{\pm 1.8}$ $65.93_{\pm.72}$ $66.7_{\pm 1.6}$ $60.8_{\pm 1.1}$ $71.5_{\pm 1.9}$ $93.72_{\pm.26}$ $\mathbf{74.72}_{\pm.14}$ $56.2_{\pm 2.0}$ $\mathbf{80.1}_{\pm 1.1}$ $\mathbf{69.7}_{\pm 2.0}$ $\mathbf{61.8}_{\pm 1.2}$
@@
C2BM $71.4_{\pm 1.7}$ $65.33_{\pm 1.1}$ $66.4_{\pm 1.5}$ $62.5_{\pm 1.4}$ $74.1_{\pm 1.8}$ $94.18_{\pm.03}$ $\mathbf{74.73}_{\pm.41}$ $58.1_{\pm 1.1}$ $\mathbf{80.5}_{\pm 0.7}$ $\mathbf{70.8}_{\pm 1.7}$ $\mathbf{60.5}_{\pm 1.4}$
@@
Table 2: Structural Hamming distance (App. F.1) and number of mistaken edges between true and learned DAG. The total number of edges is in parentheses.
@@
C2BM captures a rich causal structure that aligns well with real-world dependencies. We quantitatively assess this alignment by comparing the learned and true causal graphs in synthetic datasets. Tab. 2 reports two metrics: a structural Hamming distance (detailed in App. F.1) and the number of incorrect edges, computed after causal discovery (CD) and refinement via LLM queries. Metrics for the simplistic graphs from CBMs (all concepts are treated as mutually independent and direct causes of the task) are reported for reference. Results indicate that integrating CD with background knowledge produces a causal graph that is more accurately aligned with the true structure. Notably, on the Sachs dataset, the integration of background knowledge enables to correctly identify 10 additional edges w.r.t. CD alone. A detailed ablation study on causal graph discovery methods is provided in App. G.4. To further validate the quality of the learned causal graph, App. G.2 demonstrates that C2BM achieves comparable task accuracy using either the learned or the true graph. When considered together, the results in Tab. 1-2 highlight C2BM’s ability to improve on causal reliability without compromising expressivity and performance.
@@ 5.2 Ground-truth interventions
Figure 4: Label accuracy (%) on downstream variables (task included) after intervening on concepts up to progressively deeper levels in the graph hierarchy. Summit plots show the difference of C2BM’s accuracy w.r.t. the best-performing baseline. Uncertainties represent $2$ sample mean $\sigma$ (5 seeds).
@@ 5.3 Debiasing
We hypothesize that a real-world-aligned causal bottleneck can reduce reliance on spurious correlations. To test this, we use cMNIST, where digit Color is correlated with Parity during training (all the odd digits are green). At test time, the correlation among Color and Parity is reversed (all the even digits are green), introducing a distribution shift that challenges generalization. The neural encoders in all models still struggle with out-of-distribution (OOD) generalization.
@@
Therefore, as expected, all models including C2BM, fail to extrapolate correctly, capturing the artificial correlation. However, their enforced reasoning is different. All baselines retain the concept-task connections, perpetuating the color-parity shortcut. In contrast, C2BM detects the color-parity edge through causal discovery but correctly removes it via graph refinement with the LLM block. This is reflected in large differences in performance after concept interventions, which alleviate (or even remove) the impact of the encoder.
@@ 5.4 Fairness
We measure the Causal Concept Effect (CaCE) (Goyal et al., 2019) of Attractive on Should be Hired before and after blocking the only path between them, i.e., performing an intervention on Qualified. Tab. 3 shows that C2BM is the only model able to successfully removes the influence, achieving a post-intervention CaCE of 0.0%. This difference stems from the model architectures: in CBM, CEM, and SCBM, all concepts are directly connected to the task, meaning interventions on one concept cannot block the influence of others. In contrast, C2BM enforces a structured causal bottleneck allowing for interventions to fully override the effects of parent nodes and block information propagation through the intervened node. This highlights C2BM’s ability to enforce causal fairness by eliminating biased pathways.
@@ 6 Conclusions
C2BM has the potential to significantly narrow the hypothesis space in complex scientific domains where even a panel of human experts might struggle to identify or exclude plausible hypotheses worth testing. For instance, constructing the hypothesis space to design clinical trials accounting for the influence of environmental conditions on gastrointestinal biochemistry requires deep interdisciplinary knowledge—not only in environmental science and biochemistry, but also in genetics, microbiology, nutrition science, and epidemiology, among others. In such settings, C2BM may integrate human scientific knowledge across these diverse fields to construct a comprehensive causal graph, thereby supporting experts in systematically excluding hypotheses that are inconsistent with the integrated body of evidence. This can help accelerate interdisciplinary scientific discovery, while reducing experts’ cognitive burden.
@@
Future directions include a deeper investigation of OOD generalization with C2BMs, an extensive exploration of their role in causal inference (e.g, counterfactual queries, see App. A), and the identification of optimal intervention policies. Moreover, incorporating PAGs (Zhang, 2008) would enable modeling of hidden confounders.
Software and data
Python code and instructions for reproducing results across all datasets and methods are available at https://github.com/gdefe/causally-reliable-cbm. All datasets are open-source.
@@ Acknowledgments
This work was supported by the Hasler Foundation under the Project ID: 2024-05-15-70. AC and JS acknowledge support from FFF der universität Liechtenstein grant _lbs_24_08. SS acknowledges support from the Swiss National Science Foundation project PROSELF (No. 205121_197242). PB acknowledges support from the Swiss National Science Foundation Postdoctoral Fellowships IMAGINE (No. TMPFP2_224226). AT acknowledges the support by the Hasler Foundation grant Malescamo (No. 22050), and the Horizon Europe grant Automotif (No. 101147693).

### v3

@@ 5 Experimental evaluations
The considered datasets include both synthetic and real-world benchmarks. As synthetic datasets, we sample $10^{4}$ points from each of the five following discrete Bayesian networks available from the bnlearn repository (scutari2010): Asia (asia), Sachs (sachs2005causal), Insurance (binder1997), Alarm (alarm), and Hailfinder (abramson1996). We include cMNIST, a variant of the original dataset (lecun2010) in which the image data are colored according to custom rules. Additionally, we consider three real-world datasets: CelebA (liu2015), a facial recognition dataset labeled with different binary facial attributes; CUBC, a custom version of the original bird image dataset (he2019fine) from which we select a subset of concepts and define new ones to introduce deeper causal relationships; Siim-Pneumothorax (you2023), containing chest X-ray images annotated with a single label indicating the presence of pneumothorax, without additional concepts or their annotations. To generate them, we follow the label-free approach outlined in Sec. 4.2.1. Exhaustive details on all datasets are given in App. E.
The performance of the proposed pipeline is investigated alongside an opaque neural baseline predicting the task variable only (OpaqNN) and established state-of-the-art (SOTA) concept-based architectures, namely: CBM (koh2020), with linear and non-linear decoder; CEM (zarlenga2022concept); and SCBM (vandenhirtz2024stochastic). Hyperparameters have been selected via an independent search for each dataset–model pair based on performance on the validation set. Further details on each model’s architecture and hyperparameters can be found in App. F. Note that all baselines, except for OpaqNN, provide concept-based explanations for their predictions and allow concept interventions at test-time. This excludes architectures such as Self-Explainable Neural Networks (alvarez2018towards) and Concept Whitening (chen2020concept) as they do not offer a clear mechanism for intervening on their concept bottlenecks. We also excluded other CBM baselines such as Probabilistic CBMs (kim2023probabilistic), Post-hoc CBMs (yuksekgonul2022post), Label-free CBMs (oikarinen2023; yang2023language), as all of them share the same limitation of vanilla CBMs and CEMs: the causal graph is fixed and bipartite. Python code for reproducing all experiments is provided alongside the submission as supplementary material.
@@ 5.1 Task accuracy and causal reliability
Table 1: Task accuracy (%). Task concepts are as follows: dysp (Asia), Akt (Sachs), PropCost (Insurance), BP (Alarm), R5Fcst (Hailfinder), parity (cMNIST), mouth slightly open (CelebA), Survival (CUBC), pneumothorax (Pneumoth.). refers to reduced concept bottlenecks. Methods matching the performance of OpaQNN and showing a significant improvement over the other considered methods are highlighted in bold. Uncertainties represent 22 sample mean $\sigma$ across 5 runs.
@@
CBM+mlp $71.2_{\pm 1.4}$ $5.68_{\pm.84} 66$$6.7_{\pm 1.4}$ $62.3_{\pm 1.8}$ $70.9_{\pm 2.2}$ $93.55_{\pm.31}$ $71.27_{\pm.20}$ $56.2_{\pm 1.9}$ $76.7_{\pm 0.6}$ $58.6_{\pm 2.7}$ $52.8_{\pm 1.2}$
CEM $71.1_{\pm 1.8}$ $5.93_{\pm.72} 66$$6.7_{\pm 1.6}$ $60.8_{\pm 1.1}$ $71.5_{\pm 1.9}$ $93.72_{\pm.26}$ $\mathbf{74.72}_{\pm.14}$ $56.2_{\pm 2.0}$ $\mathbf{80.1}_{\pm 1.1}$ $\mathbf{69.7}_{\pm 2.0}$ $\mathbf{61.8}_{\pm 1.2}$
@@
C2BM $71.4_{\pm 1.7}$ $5.33_{\pm 1.1} 66$$6.4_{\pm 1.5}$ $62.5_{\pm 1.4}$ $74.1_{\pm 1.8}$ $94.18_{\pm.03}$ $\mathbf{74.73}_{\pm.41}$ $58.1_{\pm 1.1}$ $\mathbf{80.5}_{\pm 0.7}$ $\mathbf{70.8}_{\pm 1.7}$ $\mathbf{60.5}_{\pm 1.4}$
@@
C2BM captures a rich causal structure that aligns well with real-world dependencies. We quantitatively assess this alignment by comparing the learned and true causal graphs in synthetic datasets. Tab. 2 reports two metrics: a structural Hamming distance (detailed in App. F.1) and the number of incorrect edges, computed after causal discovery (CD) and refinement via LLM queries. Metrics for the simplistic graphs from CBMs (all concepts are treated as mutually independent and direct causes of the task) are reported for reference. Results indicate that integrating CD with background knowledge produces a causal graph that is more accurately aligned with the true structure. Notably, on the Sachs dataset,
Table 2: Structural Hamming distance (App. F.1) and number of mistaken edges between true and learned DAG. Reliability of standard flat CBMs is reported for reference. The total number of edges is in parentheses.
@@
the integration of background knowledge enables to correctly identify 10 additional edges w.r.t. CD alone. Detailed ablation studies on causal graph discovery methods and LLM types are provided in App. G.3-G.4-G.5. To further validate the quality of the learned causal graph, App. G.2 demonstrates that C2BM achieves comparable task accuracy using either the learned or the true graph. When considered together, the results in Tab. 1-2 highlight C2BM’s ability to improve on causal reliability without compromising expressivity and performance.
@@ 5.2 Ground-truth interventions
Figure 4: Label accuracy (%) on downstream variables (task included) after intervening on concepts up to progressively deeper levels in the graph hierarchy. Summit plots show the difference of C2BM’s accuracy w.r.t. the best-performing baseline. Uncertainties represent 22 sample mean $\sigma$ across 5 runs.
@@ 5.3 Debiasing
We hypothesize that a real-world-aligned causal bottleneck can reduce reliance on spurious correlations. To test this, we use cMNIST, where digit Color is correlated with Parity during training (all the odd digits are green). At test time, the correlation among Color and Parity is reversed (all the even digits are green), introducing a distribution shift that challenges generalization. The neural encoders in all models still struggle with out-of-distribution (OOD) generalization. Therefore, as expected, all models including C2BM, fail to extrapolate correctly, capturing the artificial correlation. However, their enforced reasoning is different. All baselines retain the concept-task connections, perpetuating the color-parity shortcut. In contrast, C2BM detects the color-parity edge through causal discovery but correctly removes it via graph refinement with the LLM block. This is reflected in large differences in performance after concept interventions, which alleviate (or even remove) the impact of the encoder.
@@ 5.4 Fairness
We measure the Causal Concept Effect (CaCE) (goyal2019) of Attractive on Should be Hired before and after blocking the only path between them, i.e., performing an intervention on Qualified. Tab. 3 shows that C2BM is the only model able to successfully remove the influence, achieving a post-intervention CaCE of 0.0%. This difference stems from the model architectures: in CBM, CEM, and SCBM, all concepts are directly connected to the task, meaning interventions on one concept cannot block the influence of others. In contrast, C2BM enforces a structured causal bottleneck allowing for interventions to fully override the effects of parent nodes and block information propagation through the intervened node. This highlights C2BM’s ability to enforce causal fairness by eliminating biased pathways.
@@ 6 Conclusions
We speculate C2BM has the potential to significantly narrow the hypothesis space in complex scientific domains where even a panel of human experts might struggle to identify or exclude plausible hypotheses worth testing. For instance, constructing the hypothesis space to design clinical trials accounting for the influence of environmental conditions on gastrointestinal biochemistry requires deep interdisciplinary knowledge—not only in environmental science and biochemistry, but also in genetics, microbiology, nutrition science, and epidemiology, among others. In such settings, C2BM may integrate human scientific knowledge across these diverse fields to construct a comprehensive causal graph, thereby supporting experts in systematically excluding hypotheses that are inconsistent with the integrated body of evidence. This can help accelerate interdisciplinary scientific discovery, while reducing experts’ cognitive burden.
@@
Future directions include a deeper investigation of OOD generalization with C2BMs, an extensive exploration of their role in causal inference (e.g, counterfactual queries extending , see App. A), and the identification of optimal intervention policies. Moreover, incorporating PAGs (Zhang2008) would enable modeling of hidden confounders.
@@ Acknowledgments
This work is supported by the Swiss National Science Foundation (SNSF) through the grant 205121_197242 for the project “PROSELF: Semi-automated Self-Tracking Systems to Improve Personal Productivity” and the Hasler Foundation under the Project ID: 2024-05-15-70. AC and JS acknowledge support from FFF of the University of Liechtenstein grant lbs_24_08. PB acknowledges support from the Swiss National Science Foundation Postdoctoral Fellowships IMAGINE (No. 224226) and has received funding from the Research Foundation Flanders (FWO, G033625N). AT acknowledges the support by the Hasler Foundation grant Malescamo (No. 22050), and the Horizon Europe grant Automotif (No. 101147693).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims of our paper have been highlighted in bold in the introduction (Sec. 1) and summarized in the abstract. Sec. 5 provide experimental evidence that support the claims presented in the abstract and introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our work are clearly stated and discussed in the Conclusions (Sec. 6).
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Although the main contribution of our work is not theoretical, we also included a simple proof regarding the expressivity of our model in App. D.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide extensive details for reproducing the experiments in Appendices B, C, C.2, E, and F. Moreover, we uploaded a .zip file containing our code.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use freely available datasets and provide instructions on how to download and preprocess them in App. E. Moreover, we have uploaded a .zip file containing our code. YAML configurations are available within the code to reproduce experiments.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Details on dataset splitting can be found in App. E, while information on training is provided in App. F.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All relevant experiments are executed with 5 random seeds. The reported results are averaged, and uncertainty is expressed for all experimental results as 2 standard errors of the sample mean. This is stated in the main results’ captions.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Information about the computational resources is provided in the App. F.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read the NeurIPS Code of Ethics and ensured that our paper conforms to them. Specifically, our experiments do not include human subjects and the content of our paper does not contain personally identifiable information.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The broader impact of our paper is discussed in the Conclusions (Sec. 6).
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: In our paper, we propose a methodological extension to a class of models that are predominantly used in scientific research. All datasets constitute standard choices in the community.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite the original creators or owners of every asset used in the paper, for example, see the references for the datasets in App. E. All our employed datasets and baselines are open-source. Datasets licenses are as follows: bnlearn datasets (CC-BY-SA), MNIST (CC BY-SA), CUB (CC0: Public Domain), CelebA (Creative Commons NonCommercial license), Pneumothorax (CC BY 4.0). As for the baselines: we implemented CEM and CBM from scratch and provide code alongside the submission. For SCBM we use the implementation available at https://github.com/mvandenhi/SCBM/tree/main.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We propose a new predictive algorithm and a pipeline to instantiate it. Detailed descriptions of how they work are provided throughout the main text and in the appendices. Moreover, we have uploaded a .zip file containing our code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does include experiments involving human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We used LLMs just for editing purposes and grammar checks. Notably, in our proposed framework, an LLM is employed as a secondary element to assist a causal discovery algorithm for causal discovery refinement.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.